OpenStudy (anonymous):

Robert purchased $2,000 worth of U.S. savings bonds. If bonds are sold in $50 or $100 denominations only, which of the following CANNOT be the number of U.S. savings bonds that Robert purchased?

OpenStudy (sasogeek):

...?

OpenStudy (anonymous):

I got 27

OpenStudy (anonymous):

no its 50 my bad

OpenStudy (sasogeek):

you said, "...which of the following....", but there's no options? :/

OpenStudy (anonymous):

29 27 30 40 50

OpenStudy (anonymous):

there they r^^^^

OpenStudy (sasogeek):

how would u solve for 27 and 29?

OpenStudy (anonymous):

50x+100y=2000 x+2y=40 neither can be 29 neither can be 27 if x is 30 then y is 5 if x is 40 then he only bought $50 bonds neither can be 50 dang! im confused now O.o

OpenStudy (sasogeek):

if it's 40, then he bought only $50 bonds, hence $50*40=$2000 if it's 30, then he bought 10 $100 bonds, and 20 $50 bonds, hence 10*$100 +20*$50=$1000+$1000=$2000 if it's 50, it's not possible, cos no matter the ratio, he will buy more than $2000 worth of bonds i don't know how to go about 27 and 29

OpenStudy (sasogeek):

are you allowed to make multiple choices of answers or just 1?

OpenStudy (anonymous):

more than one

OpenStudy (anonymous):

so its 27 and 50?

OpenStudy (anonymous):

and 29?

OpenStudy (sasogeek):

that's what i would go for, i'm convinced those are the answers :) 27,29, and 50

OpenStudy (anonymous):

Thanks so much saso!!!!!

OpenStudy (sasogeek):

anytime :)